# JAFAR: Jack up Any Feature at Any Resolution

### v1

@@ 5 Conclusion
We introduce JAFAR, a lightweight, attention-based feature upsampler designed with a simple training objective. It can upscale features from any foundation vision encoder to arbitrary output resolutions—without requiring supervision at the original image size or annotations from downstream tasks. Although task-agnostic, JAFAR outperforms prior state-of-the-art upsamplers across a variety of downstream tasks, despite not being trained specifically for them. This work lays the groundwork for a unified feature upsampler that could enable significantly more efficient architectures for dense vision tasks. Currently, the method requires training a separate upsampler for each backbone. Future work will focus on making JAFAR backbone-independent at inference time and on further reducing feature-level artifacts to produce sharper outputs.
@@
Appendix A Class Activation Maps Evaluation
@@
[…]
@@
[…]
@@
[…]
@@
where $\odot$ denotes element-wise multiplication. This condition implies that the CAM produced from the masked image should be identical to the original CAM, ensuring that the explanation is stable. Following the approach in [44], we use the Pearson Correlation Coefficient between the original CAM and the CAM obtained after masking:
@@
where Cov denotes the covariance and $\sigma$ the standard deviation of each CAM. Since the Pearson Correlation Coefficient ranges from $-1$ to $1$, we normalize it to the range $[0,1]$ and express it as a percentage for interpretability. A coherency score of 100% indicates that the attribution method is fully invariant to input perturbations guided by its own explanations.
@@
Appendix B Additional Details on Baselines
  • Large Image: For the Large Image baseline, we upsampled the original image via bilinear upsampling and use it as input to the foundation vision encoder. During evaluation on downstream tasks (see Tabs. 1 and 2), we upsample the input to the maximum ratio that fits in memory (i.e., ×$\times 8$), and subsequently apply bilinear upsampling to the resulting feature map to match the target output resolution. Due to the high computational cost and training time, we omit results on the COCO dataset. For efficiency, on Open-Vocabulary segmentation we limit upsampling to a ×$\times 2$ ratio in Tab. 3.
  • Strided: To obtain higher-resolution feature maps, we modify the stride of the ViT backbone to produce more patches. While the stride typically equals the patch size (e.g., 14 in DINOv2), we reduce the former to 6 in our experiments corresponding to a ×$\times 2.3$ upsampling. We then apply bilinear upsampling to the resulting feature map to reach the desired output resolution.
  • CARAFE [22]: For CARAFE, we use the CARAFEPack module from MMCV [51], stacking four upsampling stages with a ×$\times 2$ ratio each, resulting in a final feature map upsampled by a factor of ×$\times 16$.
  • SAPA [23]: We adopt the default implementation from the official SAPA repository, stacking four SAPA upsampling modules, each with an upsampling factor of 2. This results in a final feature map with a total upsampling factor of ×$\times 16$.
  • DySample [24]: We adopt the default implementation from the official Dysample repository, stacking four Dysample upsampling modules, each with an upsampling factor of 2. This results in a final feature map with a total upsampling factor of ×$\times 16$.
  • ReSFU [25]: We use the default implementation from the official ReSFU repository, performing a direct upsampling to the target output resolution.
  • FeatUp [28]: For FeatUp, we use the scalable JBU variant from the official FeatUp respository stacking 4 upsampling modules to achieve a total upsampling factor of ×$\times 16$. We re-train FeatUp using the provided training scripts. While the original paper trains FeatUp on the COCO dataset for 2,000 steps with a batch size of 4 and evaluates on the same dataset, we train it on the ImageNet training set for 50,000 steps (25×25$\times$25 × more) using the same batch size, ensuring a fair comparison across methods.
  • LiFT [27]: For LiFT, we slightly adapt the official implementation by resizing the intermediate representations after the downsampling module to ensure compatibility with backbones using a patch size of 14 (i.e., a downsampling factor of 14). In the official code, LiFT performs a ×$\times 2$ upsampling and then relies on a bilinear interpolation to upsample the features to the target output resolution.
We summarize in Tab. 6 the differences between upsamplers.
@@ Appendix B Additional Details on Baselines
Appendix C Additional Visualizations
We provide in the following subsections additional comparison visualizations for upsampled feature maps Sec. C.1, class activation maps predictions Sec. C.2, depth estimation Sec. C.3 and semantic segmentation Sec. C.4.
C.1 Feature Visualization
@@
Fig. 5 shows additional PCA visualizations of upsampled feature maps. Starting from $32\times 32$ feature maps extracted using a DINOv2-S/14 backbone, each is upsampled to $448\times 448$ using different baseline methods. These baselines—whether training-free, task-dependent, or task-agnostic—tend to introduce varying degrees of blurriness and visual artifacts. In contrast, JAFAR, while remaining task-agnostic, produces sharp, content-aware features with minimal artifacts.
C.2 Class Activation Maps
We present additional Grad-CAM visualizations based on ViT-B/16 features from the ImageNet validation set in Fig. 6. Except for the “Low-Res” column, where features remain at their original 14×14 resolution, all feature maps are upsampled to 224×224 before Grad-CAM extraction. The explainability maps generated by our upsampling approach are noticeably sharper and more accurate, exhibiting fewer artifacts compared to those from alternative methods. Notably, CARAFE and LiFT fail to produce meaningful explanations in this setting, suggesting that the training of these methods does not transfer effectively to these ViT-based features.
@@ C.2 Class Activation Maps
C.3 Depth Estimation
Fig. 7 presents additional examples of linear probe transfer learning for depth estimation on the COCO-Stuff dataset. Feature maps of size $32\times 32$, extracted from a DINOv2-S/14 backbone, are upsampled to $448\times 448$ using the various baseline methods. A linear probe is then trained on these features to predict depth, using supervision from a Depth-AnythingV2 model. The results demonstrate that both FeatUp variants produce high-quality features well-suited for transfer learning in depth estimation tasks.
@@ C.3 Depth Estimation
C.4 Semantic Segmentation
Fig. 8 presents examples of linear probe transfer learning for semantic segmentation on the COCO-Stuff dataset. Feature maps of size $32\times 32$, extracted from a DINOv2-S/14 backbone, are upsampled to $448\times 448$ using the various baseline methods. JAFAR produces more coherent segmentation results, offering improved delineation of both object boundaries and background regions.
@@ C.4 Semantic Segmentation
Appendix D Additionnal Comparison With Task-Agnostic Baselines
D.1 FeatUp & LiFT
In Tab. 1, we reported results obtained by training FeatUp and LiFT within our own codebase. To complement this evaluation, we present an additional comparison in Tabs. 7 and 8 using the official released checkpoints—FeatUp on DINOv2 ViT-S/14 and LiFT on DINO ViT-S/16, respectively.
@@ D.1 FeatUp & LiFT
Model 224 448 ADE20K Cityscapes VOC ADE20K Cityscapes VOC Bilinear 24.50 38.18 64.10 28.19 49.35 70.64 FeatUp 29.26 42.84 71.90 32.87 53.17 77.57 JAFAR 30.04 48.52 75.36 32.29 56.45 79.06
@@
Model 224 448 ADE20K Cityscapes VOC ADE20K Cityscapes VOC Bilinear 15.89 32.52 32.15 16.44 36.56 33.13 LiFT 16.97 36.31 35.30 16.98 40.01 35.07 LiFT-iterative 15.62 37.68 31.25 14.96 39.97 29.43 JAFAR 21.08 39.93 47.36 21.44 43.93 48.18
Across all datasets and output resolutions—with the exception of ADE20K at 448 for FeatUp—JAFAR consistently delivers significant improvements in semantic segmentation performance. We also evaluate a LiFT-iterative baseline, which stacks two LiFT ×$\times 2$ upsamplers as described in the original LiFT paper. However, this iterative approach does not outperform the simpler method of applying a single LiFT ×$\times 2$ upsampler followed by bilinear interpolation.
D.2 LoftUp
We present an additional comparison with LoftUp [53], a recent baseline which relies on segmentation masks from SAM [54] during training and employs a two-stage pipeline that includes a self-distillation phase. For a fair evaluation, we tested JAFAR within LoftUp’s official codebase using a DINOv2 ViT-S/14 backbone.
@@ D.2 LoftUp
As shown in Tab. 9, JAFAR outperforms LoftUp at lower upsampling resolutions (56 and 112) and delivers comparable performance at higher resolutions (224 and 448). Different from LoftUp, JAFAR uses a simpler and more efficient single-stage strategy: it operates entirely at low resolution and does not rely on external annotations. Nevertheless, the self-distillation mechanism introduced in LoftUp is complementary to our approach and can be seamlessly integrated into JAFAR’s pipeline. To demonstrate this, we implemented a similar distillation objective and report the results as JAFAR + distillation in Tab. 9. While the gains are minimal at lower resolutions, this enhancement provides a clear boost at higher resolutions (224 and 448). Lastly, JAFAR is considerably more lightweight, with just 0.7 million parameters compared to 4.3 million in LoftUp.
Appendix E Inference Time
In Tab. 10, we compare the inference times of various methods using a batch size of 1 and input images of resolution 448, across multiple output resolutions. The experiments are conducted on a single A100 GPU.
Table 10: Parameter Count and Inference Time Comparison (ms).
@@ Appendix E Inference Time
Additionally, Tab. 11 reports the memory usage of each model during both forward and backward passes, evaluated across multiple output resolutions using a fixed batch size of 1 and input resolution of 448.
Table 11: Memory usage (in GB) for different models and resolutions.

### v2

@@ 5 Conclusion
We introduce JAFAR, a lightweight, attention-based feature upsampler designed with a simple training objective. It can upscale features from any foundation vision encoder to arbitrary output resolutions, without requiring supervision at the original image size or annotations from downstream tasks. Although task-agnostic, JAFAR outperforms prior state-of-the-art upsamplers across a variety of downstream tasks, despite not being trained specifically for them. This work lays the groundwork for a unified feature upsampler that could enable significantly more efficient architectures for dense vision tasks. Currently, the method requires training a separate upsampler for each backbone. Future work will focus on making JAFAR backbone-independent at inference time and on further reducing feature-level artifacts to produce sharper outputs.
6 Acknowledgment
This work has been supported by chair VISA DEEP (ANR-20-CHIA-0022) Cluster PostGenAI@Paris (ANR-23-IACL-0007, FRANCE 2030). This work has been supported by PEPR Sharp (ANR-23-PEIA-0008, FRANCE 2030).
JAFAR: Jack up Any Feature at Any Resolution
@@ 6 Acknowledgment
Appendix A Additional Details
A.1 Evaluation
@@
$\textbf{A.D}=\frac{1}{N}\underset{i=1}{\overset{N}{\sum}}\;\frac{max(0,Y_{i}^{c}-O_{i}^{c})}{Y_{i}^{c}}\cdot 100,$ (4)
@@
$\textbf{A.I}=100\cdot\underset{i=1}{\overset{N}{\sum}}\;\frac{\mathbbm{1}_{Y_{i}^{c}<O_{i}^{c}}}{N},$ (5)
@@
$\textbf{A.G}=\frac{1}{N}\underset{i=1}{\overset{N}{\sum}}\;\frac{max(0,O_{i}^{c}-Y_{i}^{c})}{1-Y_{i}^{c}}\cdot 100,$ (6)
@@
where $\odot$ denotes element-wise multiplication. This condition implies that the CAM produced from the masked image should be identical to the original CAM, ensuring that the explanation is stable. Following the approach in [adcc], we use the Pearson Correlation Coefficient between the original CAM and the CAM obtained after masking:
@@
where Cov denotes the covariance and $\sigma$ the standard deviation of each CAM. Since the Pearson Correlation Coefficient ranges from 1-1 to $1$, we normalize it to the range $[0,1]$ and express it as a percentage for interpretability. A coherency score of 100% indicates that the attribution method is fully invariant to input perturbations guided by its own explanations.
@@
A.2 Baselines
  • Large Image: For the Large Image baseline, we upsampled the original image via bilinear upsampling and use it as input to the foundation vision encoder. During evaluation on downstream tasks (see Tabs.˜1 and 2), we upsample the input to the maximum ratio that fits in memory (i.e., $\times 8$), and subsequently apply bilinear upsampling to the resulting feature map to match the target output resolution. Due to the high computational cost and training time, we omit results on the COCO dataset. For efficiency, on Open-Vocabulary segmentation we limit upsampling to a $\times 2$ ratio in Tab.˜3.
  • Strided: To obtain higher-resolution feature maps, we modify the stride of the ViT backbone to produce more patches. While the stride typically equals the patch size (e.g., 14 in DINOv2), we reduce the former to 6 in our experiments corresponding to a $\times 2.3$ upsampling. We then apply bilinear upsampling to the resulting feature map to reach the desired output resolution.
  • CARAFE [carafe]: For CARAFE, we use the CARAFEPack module from MMCV [mmcv], stacking four upsampling stages with a $\times 2$ ratio each, resulting in a final feature map upsampled by a factor of $\times 16$.
  • SAPA [sapa]: We adopt the default implementation from the official SAPA repository, stacking four SAPA upsampling modules, each with an upsampling factor of 2. This results in a final feature map with a total upsampling factor of $\times 16$.
  • DySample [dysample]: We adopt the default implementation from the official Dysample repository, stacking four Dysample upsampling modules, each with an upsampling factor of 2. This results in a final feature map with a total upsampling factor of $\times 16$.
  • ReSFU [resfu]: We use the default implementation from the official ReSFU repository, performing a direct upsampling to the target output resolution.
  • FeatUp [featup]: For FeatUp, we use the scalable JBU variant from the official FeatUp respository stacking 4 upsampling modules to achieve a total upsampling factor of $\times 16$. We re-train FeatUp using the provided training scripts. While the original paper trains FeatUp on the COCO dataset for 2,000 steps with a batch size of 4 and evaluates on the same dataset, we train it on the ImageNet training set for 50,000 steps ($25\times$ more) using the same batch size, ensuring a fair comparison across methods.
  • LiFT [lift]: For LiFT, we slightly adapt the official implementation by resizing the intermediate representations after the downsampling module to ensure compatibility with backbones using a patch size of 14 (i.e., a downsampling factor of 14). In the official code, LiFT performs a $\times 2$ upsampling and then relies on a bilinear interpolation to upsample the features to the target output resolution.
We summarize in Tab.˜6 the differences between upsamplers.
@@ A.2 Baselines
Appendix B Additional Visualizations
We provide in the following subsections additional comparison visualizations for upsampled feature maps Sec.˜B.1, class activation maps predictions Sec.˜B.2, depth estimation Sec.˜B.3 and semantic segmentation Sec.˜B.4.
B.1 Feature Visualization
@@
Fig.˜5 shows additional PCA visualizations of upsampled feature maps. Starting from $32\times 32$ feature maps extracted using a DINOv2-S/14 backbone, each is upsampled to $448\times 448$ using different baseline methods. These baselines—whether training-free, task-dependent, or task-agnostic—tend to introduce varying degrees of blurriness and visual artifacts. In contrast, JAFAR, while remaining task-agnostic, produces sharp, content-aware features with minimal artifacts.
B.2 Class Activation Maps
We present additional Grad-CAM visualizations based on ViT-B/16 features from the ImageNet validation set in Fig.˜6. Except for the “Low-Res” column, where features remain at their original 14×14 resolution, all feature maps are upsampled to 224×224 before Grad-CAM extraction. The explainability maps generated by our upsampling approach are noticeably sharper and more accurate, exhibiting fewer artifacts compared to those from alternative methods. Notably, CARAFE and LiFT fail to produce meaningful explanations in this setting, suggesting that the training of these methods does not transfer effectively to these ViT-based features.
@@ B.2 Class Activation Maps
B.3 Depth Estimation
Fig.˜7 presents additional examples of linear probe transfer learning for depth estimation on the COCO-Stuff dataset. Feature maps of size $32\times 32$, extracted from a DINOv2-S/14 backbone, are upsampled to $448\times 448$ using the various baseline methods. A linear probe is then trained on these features to predict depth, using supervision from a Depth-AnythingV2 model. The results demonstrate that both FeatUp variants produce high-quality features well-suited for transfer learning in depth estimation tasks.
@@ B.3 Depth Estimation
B.4 Semantic Segmentation
Fig.˜8 presents examples of linear probe transfer learning for semantic segmentation on the COCO-Stuff dataset. Feature maps of size $32\times 32$, extracted from a DINOv2-S/14 backbone, are upsampled to $448\times 448$ using the various baseline methods. JAFAR produces more coherent segmentation results, offering improved delineation of both object boundaries and background regions.
@@ B.4 Semantic Segmentation
B.5 Attention Maps Visualization
To illustrate the behavior of the upsampling module, we visualize attention maps in Fig.˜9. For each example, a query location is selected, and its attention distribution over the low-resolution keys is shown. The maps reveal that attention tends to concentrate on the region semantically related to the query, even though all keys are accessible. This suggests that fully global attention may not be strictly necessary, and that localized variants could provide significant runtime and memory gains. Nevertheless, we argue that the mechanism should not be overly local, since feature maps produced by vision encoders often suffer from spatial misalignments and positional artifacts. Allowing each query a broader receptive field helps counteract these inconsistencies and improves upsampling quality. Conceptually, this resembles the principle of non-local means in image denoising, where a pixel is refined using information from a wider neighborhood rather than only its immediate surroundings.
Figure 9: Attention maps from JAFAR. Each map shows the attention distribution of a query location (white point) over the low-resolution keys.
Appendix C Additionnal Comparisons With Task-Agnostic Baselines
C.1 FeatUp & LiFT
In Tab.˜1, we reported results obtained by training FeatUp and LiFT within our own codebase. To complement this evaluation, we present an additional comparison in Tabs.˜10 and 8 using the official released checkpoints, FeatUp on DINOv2 ViT-S/14 and LiFT on DINO ViT-S/16, respectively.
@@ C.1 FeatUp & LiFT
224 448 Model ADE20K Cityscapes VOC ADE20K Cityscapes VOC Bilinear 24.50 38.18 64.10 28.19 49.35 70.64 FeatUp 29.26 42.84 71.90 32.87 53.17 77.57 JAFAR 30.04 48.52 75.36 32.29 56.45 79.06
@@
224 448 Model ADE20K Cityscapes VOC ADE20K Cityscapes VOC Bilinear 15.89 32.52 32.15 16.44 36.56 33.13 LiFT 16.97 36.31 35.30 16.98 40.01 35.07 LiFT-iterative 15.62 37.68 31.25 14.96 39.97 29.43 JAFAR 21.08 39.93 47.36 21.44 43.93 48.18
Across all datasets and output resolutions, with the exception of ADE20K at 448 for FeatUp, JAFAR consistently delivers significant improvements in semantic segmentation performance. We also evaluate a LiFT-iterative baseline, which stacks two LiFT $\times 2$ upsamplers as described in the original LiFT paper. However, this iterative approach does not outperform the simpler method of applying a single LiFT $\times 2$ upsampler followed by bilinear interpolation.
C.2 LoftUp
We present an additional comparison with LoftUp [loftup], a recent baseline which relies on segmentation masks from SAM [sam] during training and employs a two-stage pipeline that includes a self-distillation phase. For a fair evaluation, we tested JAFAR within LoftUp’s official codebase using a DINOv2 ViT-S/14 backbone.
@@ C.2 LoftUp
As shown in Tab.˜9, JAFAR outperforms LoftUp at lower upsampling resolutions (56 and 112) and delivers comparable performance at higher resolutions (224 and 448). Different from LoftUp, JAFAR uses a simpler and more efficient single-stage strategy: it operates entirely at low resolution and does not rely on external annotations. Nevertheless, the self-distillation mechanism introduced in LoftUp is complementary to our approach and can be seamlessly integrated into JAFAR’s pipeline. To demonstrate this, we implemented a similar distillation objective and report the results as JAFAR + distillation in Tab.˜9. While the gains are minimal at lower resolutions, this enhancement provides a clear boost at higher resolutions (224 and 448). Lastly, JAFAR is considerably more lightweight, with only 700K parameters compared to 4.3M in LoftUp.
Appendix D Additional Analysis
D.1 Upsampling Ratios
To evaluate the robustness of JAFAR upsampling across different ratios, we measure its performance on a linear probing semantic segmentation task at multiple scales, including extreme ratios ($8^{2}\rightarrow 896^{2}$). The results indicate that JAFAR consistently sustains strong performance as the upsampling ratio grows, underscoring its ability to generalize well beyond the training range.
Table 10: JAFAR robustness across upsampling scales on linear probing segmentation.
Input Resolution Dataset Upsampler $56^{2}$ $112^{2}$ $224^{2}$ $448^{2}$ $896^{2}$ $8^{2}$ Cityscapes Bilinear 31.87 31.81 31.90 31.61 31.92 JAFAR 36.54 36.04 35.71 35.77 35.54 VOC Bilinear 64.79 64.95 64.89 64.91 64.9 JAFAR 72.7 73.13 72.77 72.69 72.55 $16^{2}$ Cityscapes Bilinear 47.77 47.77 47.78 47.83 47.84 JAFAR 53.18 53.22 52.35 51.76 51.36 VOC Bilinear 75.41 75.42 75.51 75.50 75.50 JAFAR 80.57 80.76 80.58 80.36 80.24 $32^{2}$ Cityscapes Bilinear 59.56 59.53 59.41 59.54 59.15 JAFAR 59.65 61.39 61.38 60.86 60.79 VOC Bilinear 80.64 80.72 80.77 80.83 80.82 JAFAR 81.51 82.53 82.67 83.55 83.24
D.2 Positional Encoding
As shown in Tab.˜11, RoPE plays a critical role in maintaining consistent spatial alignment between queries and keys across varying spatial resolutions, with its benefits becoming more pronounced at higher resolutions.
Table 11: Ablation of RoPE on linear probing segmentation (mIoU).
Cityscapes VOC
JAFAR $56^{2}$ $112^{2}$ $224^{2}$ $448^{2}$ $56^{2}$ $112^{2}$ $224^{2}$ $448^{2}$
w/o RoPE 18.15 27.06 34.80 36.09 35.36 62.38 67.58 67.22
w/ RoPE 21.41 35.80 52.42 60.86 40.41 72.78 80.57 83.55
Appendix E Performance
We compare in Tabs.˜12 and 13 the runtime and memory usage respectively of various methods with a batch size of 1 and input resolution of 448, across multiple target resolutions. The experiments are conducted on a single A100 GPU.
Table 12: Runtime (ms) for different models and resolutions.
@@ Appendix E Performance
Table 13: Memory usage (GB) for different models and resolutions.
@@
In Tab.˜14, we report GFLOPs and training memory comparisons for the BeV segmentation setup. JAFAR delivers $6\%\!-\!15\%$ higher IoU while requiring only $3\%\!-\!10\%$ more GFLOPs than the version without upsampling, underscoring the trade-off between accuracy and computation.
Table 14: Comparison of upsampling methods on BeV segmentation pipelines.
Model Metric No Upsampling Bilinear FeatUp JAFAR LoftUp
SimpleBeV [harley2023simple] Memory (MiB) 4539 4981 5305 5453 5305
GFLOPs 594.00 595.10 595.73 612.24 621.06
mIoU 31.75 33.67 33.95 36.59
PointBeV [pointbev] Memory (MiB) 1507 1833 1961 2521 2397
GFLOPs 154.30 154.32 154.90 171.41 180.25
mIoU 34.89 36.01 35.38 37.20
BeVFormer [bevformer] Memory (MiB) 2143 2489 2639 3135 2971
GFLOPs 352.40 354.70 357.72 376.74 379.82
mIoU 33.72 34.18 34.01 36.54
Appendix F Limitations
Although our method offers significant advantages over existing upsampling approaches, it also has certain limitations. Because our approach employs a global attention mechanism, each query attends to all keys, and the number of keys increases with the resolution of the input image processed by the foundation vision encoder. As a result, the computational and memory costs can become prohibitive when dealing with large key sets (e.g., $64\times 64$ or higher). While a large receptive field is desirable, attending to every key is not strictly necessary. Alternative attention mechanisms with more localized receptive fields [swin, hassani2025generalized, pan2023slide] may achieve orders-of-magnitude efficiency improvements and represent a promising direction for future work.